17.A viable count of this culture, which represents Staphylococcus aureus, was done after 120 minutes. How many CFUs are expected if you plated 0.2 mL of a 104 dilution? (Hint: Consider the average number of cells per CFU) 18. At 65 minutes, the cells were transferred to a new medium in which the growth rate (H) was three times lower. What is the expected number of cells after 60 minutes of growth following the transfer? 19.By what factor did the population size increase between 100 and 220 minutes? 20.The culture illustrated in the graphs represents one liter of medium containing a limiting amount of glucose (0.2 grams). At time 0 the one liter culture contained 0.001g of cells. Given that stationary phase starts after 260 minutes of exponential growth, what would be the growth yield of this culture (Yg Grams of bacteria/gram of glucose)? Refer to the graphs on pages 4 and 5 to answer the questions. Note both graphs present the same data using different scales for the Y axis.
Answer #1

Average number of cells or CFU = \frac{number of cells}{dilution X volume of sample}

According to given graph, number of cells at 120 minutes = 1 x 107

dilution of sample= 10-4

Volume of sample = 0.2mL

CFU = \frac{1 X10^{7}}{0.2 X 10^{-4}}

         = 5 x 1011

Therefore, CFU obtained by plating 0.2mL of sample at 10-4 dilution = 5 x 1011 .